If the complement of an angle is 22° less than one-third of its supplement, find the measure of the angle.
Found 2 solutions by amoresroy, Gogonati:
Answer by amoresroy(361)

Let x = the measure of the angle
90 - x = the complement of the angle
180-x = the supplement of the angle
90-x = (180-x)/3 - 22
90-x = 60 - x/3 -22
90-60+22 = -x/3 + x
52 = 2x/3
x = 78
The measure of the angle is 78 degrees

Answer by Gogonati(855)

Let = x degree the unknown angle, then its complement is 90-x, and its supplement is 180-x. Now we write the equation: 90-x=(1/3)(180-x)-22, solving this equation we get:x=78 degree.
Answer: The unknown angle is 78 degree.